OpenStudy (anonymous):

Simplify please! x+10 over 10 = 18 over 12.

OpenStudy (anonymous):

first multiply 10 to both sides to get the variable by itself. You will then have: x+10 = 180/12 x+10 = 15 then subtract 10 from both sides to get x x = 5

OpenStudy (anonymous):

thank you very much! ♥